OpenStudy (anonymous):

What does x equal? x + 5/8 = 0

OpenStudy (dmndlife24):

The opposite of 5/8

sammixboo (sammixboo):

Well, first you can do 5/8, so what is 5/8?

OpenStudy (dmndlife24):

If you subtract 5/8 from both sides you will find that x equals -5/8, the opposite of 5/8. That's how I look at it :D

OpenStudy (anonymous):

Thats what I was thinking just didnt know if i was right.. thanks:)

sammixboo (sammixboo):

See you can do 5/8 which equals 0.625, so you can plug in 0.625 where 5/8 was. x + 0.625 = 0 Then you can subtract 0.625 on both sides x + 0.625 - 0.625 = 0 - 0.625 x = -0.625 -0.625 as a fraction is \(\sf \frac{-5}{8}\), so x = -0.625 or \(\sf \frac{-5}{8}\)

sammixboo (sammixboo):

Or you can do what @dmndlife24 said :P Whichever way seems easier, because I know it can get confusing to subtract fractions someimes

sammixboo (sammixboo):

If you have anymore questions, tag me \(\sf \color{red}{@sammixboo}\) or Dmnd \(\sf \color{red}{@dmndlife24}\)

OpenStudy (dmndlife24):

That's definitely true :) @sammixboo And thanks!

sammixboo (sammixboo):

No problem!